# ExGRG: Explicitly-Generated Relation Graph for Self-Supervised Representation Learning

### v1

@@ 3.3.2. Higher-Order Graph Encodings
3.3.3. Deep Clustering for Relation Graph Generation
@@
(5) ${\bm{P}}_{j,k}=\frac{\exp(\frac{1}{\tau}{\bm{H}}_{j,.}^{T}{\bm{C}}_{k,.})}{\sum_{k^{\prime}}\exp(\frac{1}{\tau}{\bm{H}}_{j,.}^{T}{\bm{C}}_{k^{\prime},.})}.$
@@
(6) $\begin{split}\mathcal{L}_{O}=\frac{1}{|S_{O}|}\sum_{(i,j)\in S_{O}}l_{i,j},\text{~{}where~{}}l_{i,j}=-\sum_{k}{\bm{Q}}_{i,k}\log{\bm{P}}_{j,k},\end{split}$
(7) and $\begin{split}\text{and }S_{O}=\{(i,j)\in\mathbb{N}^{2}~{}|~{}i=j\}\cup S_{a}.\end{split}$
@@
3.4. Multi-Source Aggregation
@@
(9) ${\bm{G}}=\sum_{{\bm{G}}^{(i)}\in S_{G}}\lambda_{i}F_{SG}({\bm{G}}^{(i)}),$
(10) where $\begin{split}\text{where }S_{G}=\{{\bm{G}}^{a},{\bm{G}}^{k},{\bm{G}}^{A},{\bm{G}}^{A^{\prime}},{\bm{G}}^{L},{\bm{G}}^{R},{\bm{G}}^{R^{\prime}},{\bm{G}}^{S},{\bm{G}}^{O}\}.\end{split}$
@@
(11) $\lambda_{i}=\Psi_{\psi}\left(F_{s}({\bm{G}}^{(i)})\right),$
@@
3.5. End-to-End Training Procedure
The transition between optimizing our modules, specifically the alteration of (i) updating $f_{\theta}$ and $g_{\phi}$, and (ii) generation of a more informative ${\bm{G}}^{(i)}\in S_{G}$ and eventually ${\bm{G}}$, can be achieved through the incorporation of an SG mechanism (Eq. 9), inspired by (Grill et al., 2020; Chen and He, 2021), and a relation matrix regularization as
(12) $\mathcal{L}_{R}=-\sum_{i,j}({\bm{G}}_{i,j})^{2}.$
@@ 3.5. End-to-End Training Procedure
Supervised GCN (Kipf and Welling, 2016a) $77.19\pm 0.12$ $86.51\pm 0.54$ $92.42\pm 0.22$ $93.03\pm 0.31$ $95.65\pm 0.16$ $81.54\pm 0.68$ $70.73\pm 0.65$ $79.16\pm 0.25$ $82.7\pm 0.00$
Node2Vec (Grover and Leskovec, 2016) $71.79\pm 0.05$ $84.39\pm 0.08$ $89.67\pm 0.12$ $85.08\pm 0.03$ $91.19\pm 0.04$ $71.08\pm 0.91$ $47.34\pm 0.84$ $66.23\pm 0.95$
DeepWalk (Perozzi et al., 2014) $74.35\pm 0.06$ $85.68\pm 0.06$ $89.44\pm 0.11$ $84.61\pm 0.22$ $91.77\pm 0.15$ $70.72\pm 0.63$ $51.39\pm 0.41$ $73.27\pm 0.86$
DW + Features (Grover and Leskovec, 2016; Lee et al., 2022) $77.21\pm 0.03$ $86.28\pm 0.07$ $90.05\pm 0.08$ $87.70\pm 0.04$ $94.90\pm 0.09$
DGI (Veličković et al., 2018) $75.35\pm 0.14$ $83.95\pm 0.47$ $91.61\pm 0.22$ $92.15\pm 0.63$ $94.51\pm 0.52$ $82.34\pm 0.71$ $71.83\pm 0.54$ $76.78\pm 0.31$
GMI (Peng et al., 2020) $74.85\pm 0.08$ $82.21\pm 0.31$ $90.68\pm 0.17$ OOM OOM $82.39\pm 0.65$ $71.72\pm 0.15$ $79.34\pm 1.04$
MVGRL (Hassani and Khasahmadi, 2020) $77.52\pm 0.08$ $87.52\pm 0.11$ $91.74\pm 0.07$ $92.11\pm 0.12$ $95.33\pm 0.03$ $83.45\pm 0.68$ $73.28\pm 0.48$ $80.09\pm 0.62$
GRACE (Zhu et al., 2020b) $77.97\pm 0.63$ $86.50\pm 0.33$ $92.46\pm 0.18$ $92.17\pm 0.04$ OOM $81.92\pm 0.89$ $71.21\pm 0.64$ $80.54\pm 0.36$
GCA (Zhu et al., 2021b) $77.94\pm 0.67$ $87.32\pm 0.50$ $92.39\pm 0.33$ $92.84\pm 0.15$ OOM $82.07\pm 0.10$ $71.33\pm 0.37$ $80.21\pm 0.39$
BGRL (Thakoor et al., 2021) $79.98\pm 0.10$ $90.34\pm 0.19$ $93.17\pm 0.30$ $93.31\pm 0.13$ $95.73\pm 0.05$ $83.83\pm 1.61$ $72.32\pm 0.89$ $86.03\pm 0.33$ $84.07\pm 0.23$
AFGRL (Lee et al., 2022) $77.62\pm 0.49$ $89.88\pm 0.33$ $93.22\pm 0.28$ $93.27\pm 0.17$ $95.69\pm 0.10$ $81.60\pm 0.54$ $71.02\pm 0.37$ $80.02\pm 0.48$
SPGCL (Wang et al., 2023) $79.01\pm 0.51$ $89.68\pm 0.19$ $92.49\pm 0.31$ $91.92\pm 0.10$ $95.12\pm 0.15$ $83.16\pm 0.13$ $71.96\pm 0.42$ $79.16\pm 0.73$
@@
Table 1. Downstream performance measured in terms of node classification accuracy’s mean and standard deviation over 20 random model initializations and dataset splits. OOM indicates Out Of Memory.
(13) $\mathcal{L}_{ETE}=\alpha\mathcal{L}_{V}+\beta\mathcal{L}_{C}+\gamma\mathcal{L}_{I^{\prime}}+\alpha_{1}\mathcal{L}_{O}+\alpha_{2}\mathcal{L}_{R},$
4. Experimental Evaluation
4.1. Experimental Setup
We undertake an extensive experimental analysis, aiming to demonstrate the efficacy of our approach and its superiority in comparison to other state-of-the-art methods. A GCN (Kipf and Welling, 2016a) as the encoder $f_{\theta}$ is utilized, with details provided in Appendix § B.2. Also, we follow the linear probing protocol, the established approach in prior graph SSL works (Thakoor et al., 2021). Further details can be found in § 3.1.1 and Appendix § B.3.
Baselines: We compare ExGRG with various graph representation learning methods, encompassing (i) supervised MLP employing raw node features and GCN (Kipf and Welling, 2016a), (ii) conventional unsupervised graph embedding works Node2Vec (Grover and Leskovec, 2016) and DeepWalk (Perozzi et al., 2014), alongside (iii) state-of-the-art contrastive and non-contrastive SSL methods, including DGI (Veličković et al., 2018), GMI (Peng et al., 2020), MVGRL (Hassani and Khasahmadi, 2020), GRACE (Zhu et al., 2020b), GCA (Zhu et al., 2021b), BGRL (Thakoor et al., 2021), and augmentation-free works AFGRL (Lee et al., 2022) and SPGCL (Wang et al., 2023).
4.2. Experimental Results
@@
4.3. Ablation Studies
@@
Table 2. Ablation studies on Amazon Computers. Models are altered concerning ExGRG (5k iters) as the reference. The metrics are outlined in Appendix § B.4.
@@ 4.4. SSL Pre-Training Analysis
Figure 2. Accuracy vs. iterations over 20 trials throughout the SSL pre-training.
@@
4.5. Hyperparameter Analysis
Figure 3. Impact of altering $D_{H}$ on ranks and downstream accuracy for Amazon Computers.
@@ A.1. Preliminary
A.2. Explicit Compositional Relation Graph
@@
A.3. Laplacian Eigenmap
(14) ${\bm{Z}}^{*}=\operatorname*{arg\,min}_{{\bm{Z}}}\operatorname{Tr}\left({\bm{Z}}^{T}{\bm{L}}{\bm{Z}}\right)\text{ s.t. }{\bm{Z}}^{T}{\bm{D}}{\bm{Z}}={\bm{I}},$
@@ A.3. Laplacian Eigenmap
A.4. Neighborhood Similarity in the Representation Space
@@
A.5. Higher-Order Graph Encodings
A.5.1. Adjacency-based Relation Graph
@@
A.5.2. PSE-based Relation Graph
@@
A.6. Deep Clustering
To obtain the codes ${\bm{Q}}$, we follow the approach outlined in (Asano et al., 2019; Caron et al., 2020). This is performed by solving the optimization problem by maximizing the similarity between representations and prototypes, with ${\bm{Q}}^{*}$ representing the optimal solution as
(15) ${\bm{Q}}^{*}=\operatorname*{arg\,max}_{{\bm{Q}}}\text{Tr}({\bm{Q}}^{T}{\bm{C}}{\bm{H}}^{T})+{\epsilon}\mathscr{H}({\bm{Q}}).$
@@ A.6. Deep Clustering
(16) $\mathcal{Q}=\left\{{\bm{Q}}\in\mathbb{R}_{+}^{N\times K}~{}|~{}{\bm{Q}}^{T}\mathbf{1}_{N}=\frac{1}{K}\mathbf{1}_{K},{\bm{Q}}\mathbf{1}_{K}=\frac{1}{N}\mathbf{1}_{N}\right\}.$
@@
(17) ${\bm{Q}}^{*}=\text{Diag}(\mathbf{u})\exp\left(\frac{{\bm{C}}{\bm{H}}^{T}}{{\epsilon}}\right)\text{Diag}(\mathbf{v}).$
A.7. End-to-End Training Procedure
A.7.1. SG Mechanism
In the presence of a gradient path through the invariance term $\mathcal{L}_{I^{\prime}}$, the optimizer incentivizes the reduction of ${\bm{G}}^{(i)}$ entries. However, it is essential to note that these specific entries within ${\bm{G}}^{(i)}$ are a refined sparse subset derived from a potentially dense $N\times N$ relation matrix. The existence of such a gradient path to update ${\bm{G}}^{(i)}$ generator parameters implies the promotion of weaker relations among instances that initially exhibit strong relations. Inspired by (Grill et al., 2020; Chen and He, 2021), to simultaneously optimize both sets of parameters, we propose an SG mechanism specifically on the ${\bm{G}}^{(i)}$ generator parameters (Eq. 9), such as the prototypes ${\bm{C}}$ for ${\bm{G}}^{O}$, through the $\mathcal{L}_{I^{\prime}}$ path. Without this SG, we counter-intuitively diminish the ${\bm{G}}^{(i)}$ entries for candidate pairs with high potential for closer representations.
A.7.2. Relation Matrix Regularization
@@
A.7.3. Overall Multi-term Loss Objective
@@ Appendix B Experimental Evaluation
B.1. Datasets
@@
B.2. SSL Encoder
We employ the GCN proposed in (Kipf and Welling, 2016a) as the SSL encoder $f_{\theta}$. Specifically, the output of the $l$-th GCN layer, ${H}^{(l)}$, is expressed as
(18) ${H}^{(l)}=\text{GCN}^{(l)}{({\bm{X}},{\bm{A}})}=\sigma(\hat{{\bm{D}}}^{-1/2}{\hat{{\bm{A}}}}{\hat{{\bm{D}}}}^{-1/2}{{\bm{X}}{\bm{W}}}^{(l)}).$
@@ B.3. Downstream Evaluation
B.4. Additional Evaluation Metrics
@@
B.5. Ablation Studies
@@
Table 4. Ablation studies on CiteSeer. Models are altered concerning ExGRG (5k iters) as the reference.
@@
Table 5. Ablation studies on Cora. Models are altered concerning ExGRG (5k iters) as the reference.
@@ B.6. Pre-Training Analysis
Figure 4. corr ${\bm{H}}$ vs. iterations throughout the SSL pre-training.
Figure 5. corr ${\bm{Z}}$ vs. iterations throughout the SSL pre-training.
@@
Figure 6. nstd ${\bm{H}}$ vs. iterations throughout the SSL pre-training.
Figure 7. std ${\bm{Z}}$ vs. iterations throughout the SSL pre-training.
@@
Figure 8. nrank ${\bm{H}}$ vs. iterations throughout the SSL pre-training.
Figure 9. nrank ${\bm{Z}}$ vs. iterations throughout the SSL pre-training.
@@
B.7. Selected Hyperparameters

### v2

@@ 3.3.2 Higher-Order Graph Encodings
3.3.3 Deep Clustering for Relation Graph Generation
@@
${\bm{P}}_{j,k}=\frac{\exp(\frac{1}{\tau}{\bm{H}}_{j,.}^{T}{\bm{C}}_{k,.})}{\sum_{k^{\prime}}\exp(\frac{1}{\tau}{\bm{H}}_{j,.}^{T}{\bm{C}}_{k^{\prime},.})}.$ (5)
@@
$\begin{split}\mathcal{L}_{O}=\frac{1}{|S_{O}|}\sum_{(i,j)\in S_{O}}l_{i,j},\text{~{}where~{}}l_{i,j}=-\sum_{k}{\bm{Q}}_{i,k}\log{\bm{P}}_{j,k},\end{split}$ (6)
and $\begin{split}\text{and }S_{O}=\{(i,j)\in\mathbb{N}^{2}~{}|~{}i=j\}\cup S_{a}.\end{split}$ (7)
@@
3.4 Multi-Source Aggregation
@@
${\bm{G}}=\sum_{{\bm{G}}^{(i)}\in S_{G}}\lambda_{i}F_{SG}({\bm{G}}^{(i)}),$ (9)
where $\begin{split}\text{where }S_{G}=\{{\bm{G}}^{a},{\bm{G}}^{k},{\bm{G}}^{A},{\bm{G}}^{A^{\prime}},{\bm{G}}^{L},{\bm{G}}^{R},{\bm{G}}^{R^{\prime}},{\bm{G}}^{S},{\bm{G}}^{O}\}.\end{split}$ (10)
@@
$\lambda_{i}=\Psi_{\psi}\left(F_{s}({\bm{G}}^{(i)})\right),$ (11)
@@
3.5 End-to-End Training Procedure
The transition between optimizing our modules, specifically the alteration of (i) updating $f_{\theta}$ and $g_{\phi}$, and (ii) generation of a more informative ${\bm{G}}^{(i)}\in S_{G}$ and eventually ${\bm{G}}$, can be achieved through the incorporation of an SG mechanism (Eq. 9), inspired by Grill et al. (2020) and Chen & He (2021), alongside a relation matrix regularization as
$\mathcal{L}_{R}=-\sum_{i,j}({\bm{G}}_{i,j})^{2}.$ (12)
@@ 3.5 End-to-End Training Procedure
$\mathcal{L}_{ETE}=\alpha\mathcal{L}_{V}+\beta\mathcal{L}_{C}+\gamma\mathcal{L}_{I^{\prime}}+\alpha_{1}\mathcal{L}_{O}+\alpha_{2}\mathcal{L}_{R},$ (13)
4 Experimental Evaluation
4.1 Experimental Setup
We undertake an extensive experimental analysis, aiming to demonstrate the efficacy of our approach and its superiority in comparison to other state-of-the-art methods. A GCN (Kipf & Welling, 2016a) as the encoder $f_{\theta}$ is utilized, with details provided in Appendix § B.2. Also, we follow the linear probing protocol, the established approach in prior graph SSL works (Thakoor et al., 2021). Further details can be found in § 3.1.1 and Appendix § B.3.
Baselines: We compare ExGRG with various graph representation learning methods, encompassing (i) supervised MLP employing raw node features and GCN (Kipf & Welling, 2016a), (ii) conventional unsupervised graph embedding works Node2Vec (Grover & Leskovec, 2016) and DeepWalk (Perozzi et al., 2014), alongside (iii) state-of-the-art contrastive and non-contrastive SSL methods, including DGI (Veličković et al., 2018), GMI (Peng et al., 2020), MVGRL (Hassani & Khasahmadi, 2020), GRACE (Zhu et al., 2020c), GCA (Zhu et al., 2021b), BGRL (Thakoor et al., 2021), and augmentation-free works AFGRL (Lee et al., 2022) and SPGCL (Wang et al., 2023).
4.2 Experimental Results
Table 1: Downstream performance measured in terms of node classification accuracy’s mean and standard deviation over 20 random model initializations and dataset splits. OOM indicates Out Of Memory.
@@ 4.2 Experimental Results
Supervised GCN $77.19\pm 0.12$ $86.51\pm 0.54$ $92.42\pm 0.22$ $93.03\pm 0.31$ $95.65\pm 0.16$ $81.54\pm 0.68$ $70.73\pm 0.65$ $79.16\pm 0.25$ $82.7\pm 0.00$
Node2Vec $71.79\pm 0.05$ $84.39\pm 0.08$ $89.67\pm 0.12$ $85.08\pm 0.03$ $91.19\pm 0.04$ $71.08\pm 0.91$ $47.34\pm 0.84$ $66.23\pm 0.95$
DeepWalk $74.35\pm 0.06$ $85.68\pm 0.06$ $89.44\pm 0.11$ $84.61\pm 0.22$ $91.77\pm 0.15$ $70.72\pm 0.63$ $51.39\pm 0.41$ $73.27\pm 0.86$
DW + Features $77.21\pm 0.03$ $86.28\pm 0.07$ $90.05\pm 0.08$ $87.70\pm 0.04$ $94.90\pm 0.09$
DGI $75.35\pm 0.14$ $83.95\pm 0.47$ $91.61\pm 0.22$ $92.15\pm 0.63$ $94.51\pm 0.52$ $82.34\pm 0.71$ $71.83\pm 0.54$ $76.78\pm 0.31$
GMI $74.85\pm 0.08$ $82.21\pm 0.31$ $90.68\pm 0.17$ OOM OOM $82.39\pm 0.65$ $71.72\pm 0.15$ $79.34\pm 1.04$
MVGRL $77.52\pm 0.08$ $87.52\pm 0.11$ $91.74\pm 0.07$ $92.11\pm 0.12$ $95.33\pm 0.03$ $83.45\pm 0.68$ $73.28\pm 0.48$ $80.09\pm 0.62$
GRACE $77.97\pm 0.63$ $86.50\pm 0.33$ $92.46\pm 0.18$ $92.17\pm 0.04$ OOM $81.92\pm 0.89$ $71.21\pm 0.64$ $80.54\pm 0.36$
GCA $77.94\pm 0.67$ $87.32\pm 0.50$ $92.39\pm 0.33$ $92.84\pm 0.15$ OOM $82.07\pm 0.10$ $71.33\pm 0.37$ $80.21\pm 0.39$
BGRL $79.98\pm 0.10$ $90.34\pm 0.19$ $93.17\pm 0.30$ $93.31\pm 0.13$ $95.73\pm 0.05$ $83.83\pm 1.61$ $72.32\pm 0.89$ $86.03\pm 0.33$ $84.07\pm 0.23$
AFGRL $77.62\pm 0.49$ $89.88\pm 0.33$ $93.22\pm 0.28$ $93.27\pm 0.17$ $95.69\pm 0.10$ $81.60\pm 0.54$ $71.02\pm 0.37$ $80.02\pm 0.48$
SPGCL $79.01\pm 0.51$ $89.68\pm 0.19$ $92.49\pm 0.31$ $91.92\pm 0.10$ $95.12\pm 0.15$ $83.16\pm 0.13$ $71.96\pm 0.42$ $79.16\pm 0.73$
@@
4.3 Ablation Studies
Table 2: Ablation studies on Amazon Computers. Models are altered concerning ExGRG (5k iters) as the reference. The metrics are outlined in Appendix § B.4.
@@ 4.4 SSL Pre-Training Analysis
Figure 2: Accuracy vs. iterations over 20 trials throughout the SSL pre-training.
@@
4.5 Hyperparameter Analysis
Figure 3: Impact of altering $D_{H}$ on ranks and downstream accuracy for Amazon Computers.
@@ A.1 Preliminary
A.2 Explicit Compositional Relation Graph
@@
A.3 Laplacian Eigenmap
${\bm{Z}}^{*}=\operatorname*{arg\,min}_{{\bm{Z}}}\operatorname{Tr}\left({\bm{Z}}^{T}{\bm{L}}{\bm{Z}}\right)\text{ s.t. }{\bm{Z}}^{T}{\bm{D}}{\bm{Z}}={\bm{I}},$ (14)
@@ A.3 Laplacian Eigenmap
A.4 Neighborhood Similarity in the Representation Space
@@
A.5 Higher-Order Graph Encodings
A.5.1 Adjacency-based Relation Graph
@@
A.5.2 PSE-based Relation Graph
@@
A.6 Deep Clustering
To obtain the codes ${\bm{Q}}$, we follow the approach outlined in Asano et al. (2019) and Caron et al. (2020). This is performed by solving the optimization problem by maximizing the similarity between representations and prototypes, with ${\bm{Q}}^{*}$ representing the optimal solution as
${\bm{Q}}^{*}=\operatorname*{arg\,max}_{{\bm{Q}}}\text{Tr}({\bm{Q}}^{T}{\bm{C}}{\bm{H}}^{T})+{\epsilon}\mathscr{H}({\bm{Q}}).$ (15)
@@ A.6 Deep Clustering
$\mathcal{Q}=\left\{{\bm{Q}}\in\mathbb{R}_{+}^{N\times K}~{}|~{}{\bm{Q}}^{T}\mathbf{1}_{N}=\frac{1}{K}\mathbf{1}_{K},{\bm{Q}}\mathbf{1}_{K}=\frac{1}{N}\mathbf{1}_{N}\right\}.$ (16)
@@
${\bm{Q}}^{*}=\text{Diag}(\mathbf{u})\exp\left(\frac{{\bm{C}}{\bm{H}}^{T}}{{\epsilon}}\right)\text{Diag}(\mathbf{v}).$ (17)
A.7 End-to-End Training Procedure
A.7.1 SG Mechanism
In the presence of a gradient path through the invariance term $\mathcal{L}_{I^{\prime}}$, the optimizer incentivizes the reduction of ${\bm{G}}^{(i)}$ entries. However, it is essential to note that these specific entries within ${\bm{G}}^{(i)}$ are a refined sparse subset derived from a potentially dense $N\times N$ relation matrix. The existence of such a gradient path to update ${\bm{G}}^{(i)}$ generator parameters implies the promotion of weaker relations among instances that initially exhibit strong relations. Inspired by Grill et al. (2020) and Chen & He (2021), to simultaneously optimize both sets of parameters, we propose an SG mechanism specifically on the ${\bm{G}}^{(i)}$ generator parameters (Eq. 9), such as the prototypes ${\bm{C}}$ for ${\bm{G}}^{O}$, through the $\mathcal{L}_{I^{\prime}}$ path. Without this SG, we counter-intuitively diminish the ${\bm{G}}^{(i)}$ entries for candidate pairs with high potential for closer representations.
A.7.2 Relation Matrix Regularization
@@
A.7.3 Overall Multi-term Loss Objective
@@ Appendix B Experimental Evaluation
B.1 Datasets
Table 3: Statistics of the employed datasets.
@@ B.1 Datasets
B.2 SSL Encoder
We employ the GCN proposed in Kipf & Welling (2016a) as the SSL encoder $f_{\theta}$. Specifically, the output of the $l$-th GCN layer, ${H}^{(l)}$, is expressed as
${H}^{(l)}=\text{GCN}^{(l)}{({\bm{X}},{\bm{A}})}=\sigma(\hat{{\bm{D}}}^{-1/2}{\hat{{\bm{A}}}}{\hat{{\bm{D}}}}^{-1/2}{{\bm{X}}{\bm{W}}}^{(l)}).$ (18)
@@ B.3 Downstream Evaluation
B.4 Additional Evaluation Metrics
@@
B.5 Ablation Studies
Table 4: Ablation studies on CiteSeer. Models are altered concerning ExGRG (5k iters) as the reference.
@@ B.5 Ablation Studies
Table 5: Ablation studies on Cora. Models are altered concerning ExGRG (5k iters) as the reference.
@@ B.6 Pre-Training Analysis
Figure 4: corr ${\bm{H}}$ vs. iterations throughout the SSL pre-training.
Figure 5: corr ${\bm{Z}}$ vs. iterations throughout the SSL pre-training.
@@
Figure 6: nstd ${\bm{H}}$ vs. iterations throughout the SSL pre-training.
Figure 7: std ${\bm{Z}}$ vs. iterations throughout the SSL pre-training.
@@
Figure 8: nrank ${\bm{H}}$ vs. iterations throughout the SSL pre-training.
Figure 9: nrank ${\bm{Z}}$ vs. iterations throughout the SSL pre-training.
@@
B.7 Selected Hyperparameters
Table 6: Selected hyperparameters utilized for each dataset.